Mathematical Methods for Social Scientists Math 195 (Sec 91), Summer 2006 Course Document

Lecturer: David Rule Office: Eckhart 16 Phone: (773) 702 6045 Email: rule@math.uchicago.edu Website: www.math.uchicago.edu/˜ rule/195/

Lectures: Monday, Wednesday & Friday, 9am-11pm, SS108 Tutorial: TBA Reader: TBA

Office hours: By appointment.

Text: Multivariable Calculus, Stewart, Chapters 11, 13, 14, 15 & 16.

Syllabus: The aim of this course is to understand the methods of multivariable calculus, with an emphasis on applications. The course covers parametric equations, vectors, the dot and cross product, functions of several variables, partial derivatives, Lagrange multipliers, and double and triple integration.

Assessment: Your grade will be determined by five in-class tests and homework. The first four tests constitute 15% of your grade, the last test 30%, and homeworks total 10%. Any student wishing to take the course for a P/F grade must request this in writing before the start of the last in-class test.

Homework: Homework should be handed in on or before the given deadline. Late homework will be given zero credit unless an extension is given by me. Extensions and the like will only be given in exceptional circumstances (for example, ill health) and should be requested before the deadline. Please write up your homework neatly and explain yourself clearly. The aim is to understand the material rather than just get the “right answer” in the end.

Miscellaneous items: You are welcome to work on problems together but must write up your solutions independently. Please refer to the University guidelines on plagiarism.

David Rule